# and Machine Learning

### v1

@@ Abstract
This paper aims to address this gap. We present a comprehensive perspective on denoisers, their structure, and desired properties. We emphasize the increasing importance of denoising and showcase its evolution into an essential building block for complex tasks in imaging, inverse problems, and machine learning. Despite its long history, the community continues to uncover unexpected and groundbreaking uses for denoising, further solidifying its place as a cornerstone of scientific and engineering practice.
@@ 1 Introduction
The conservation Property 2 guarantees that the ideal denoiser is the gradient of a scalar field. This also implies that $f(\mathbf{x},\alpha)$ is a Lipschitz map with some constant $M(\alpha)$:
@@
We naturally expect $f(\mathbf{0},\alpha)=\mathbf{0}$ for all $\alpha$; therefore, this Lipschitz condition is equivalent to $\|f(\mathbf{x},\alpha)\|\leq M(\alpha)\|\mathbf{x}\|$. A non-expansive denoiser would require that $M(\alpha)\leq 1$. In the statistics literature, such operators are called shrinking smoothers [3, 4].
@@ 3.1 Bayesian Denoisers
where $\phi(\mathbf{u})=-\log P(\mathbf{u})$ is the negative log-prior on the space of ”clean” images, and $\alpha$ is proportional to the noise variance. It would appear that the MAP denoiser does not have a closed form. However, the expression 17 is also known in the optimization literature as a proximal operator [21] when $\phi$ is convex, quasi-convex, or a difference of convex functions. It is well-known [22] that to every proximal operator $f$ there corresponds a convex scalar-valued function $\psi$ such that $f=\nabla\psi$.
@@
where $\tilde{\phi}_{\alpha}(\mathbf{x})=-\log P(\mathbf{x},\alpha)$. This is more or less identical to the form of the MAP denoiser in 18.
We graphically illustrate an example (for the scalar case) of the MMSE denoiser for $L_{1}$ loss in Fig. (5)(left). The two estimators are compared on the same plot in Fig 5 (right).
@@
Data-driven MMSE denoisers: The typical modern supervised approach to image denoising is to train a deep neural network with pairs of clean and noisy images, where the noise is often modeled as additive white and Gaussian(AWGN) [CITE]. Let’s assume we have image pairs $(\mathbf{u},\mathbf{x})\sim P(\mathbf{u},\mathbf{x})$ where $\mathbf{u}$ represents a clean image, and $\mathbf{x}$ is the noisy observation obtained by adding AWGN with a known standard deviation to $\mathbf{u}$: $\mathbf{x}=\mathbf{u}+\mathbf{e}$, where $\mathbf{e}\sim\mathcal{N}(0,\sigma^{2})$.
@@
As mentioned earlier, the MMSE denoiser is the average of all plausible clean signals given the noisy observation. This averaging can lead to a loss of details and a blurry appearance, especially when the noise level is high. Denoising Autoencoders (DAEs) are a prime example of data-driven MMSE denoisers [27]. These neural networks excel at learning robust data representations by training on noisy input and striving to reconstruct the original, clean data. This makes them not only valuable for denoising but also for tasks like data compression and feature extraction.
@@ 3.2.1 Approximation of Energy-based Denoisers
Though these facts were neither historically clarified, nor the original motivation for their development [29], denoisers of the form 31)*** have always been heuristic/empirical approximations to the MMSE. These denoisers were hugely popular and effective (e.g. [32, 11]) for decades before the more recent introduction of neural networks.
@@ 4 Denoising, the Score Function, and Generative Modeling
At first blush, it is not at all clear why this is a reasonable procedure. Yet there are a number of ways [44, 45, 43, 46, 47] to motivate this idea -perhaps none more direct than by using Tweedie’s formula introduced earlier in Eq. 20.
@@
Despite its elegance, the MMSE estimator is typically difficult to compute, or entirely inaccessible. Therefore as a proxy, often other denoisers are used, which may only be rough approximations of the MMSE (Eq. 43).
@@ 4.1 Denoising as the Engine of Diffusion Models
Denoising Diffusion Models (DDMs) [45, 43, 50, 51, 47] have become an important area of research in generative modeling. They operate by progressively corrupting training data with noise until it’s indistinguishable from random noise, then learning to systematically reverse this corruption. By training a model to iteratively denoise, it gains the ability to generate entirely new, coherent data samples from a starting point of pure noise, effectively converting noise into meaningful structures like images or other data forms (Figure  6).
@@
which we call a residual flow. As we’ve described in the previous sections, lack of access to the MMSE denoiser forces us to select a different denoiser and therefore solve only an approximate version of the desired Eq. 49.
@@ 5 Denoisers in the Context of Inverse Problems
Over the last several decades, a vast number of choices for the regularizer $\mathbf{R}(\mathbf{x},\alpha)$ have been proposed with varying degrees of success. More recently, and independently of the machine learning literature, the development of strong denoising algorithms has had a profound impact on the field of inverse problems and low-level vision, as denoisers have been shown to enable powerful priors/regularizers [55, 36].
@@
for the gradient in 57, we have the following expression for the gradient of the objective:
@@
In the particular case of RED: $\mathbf{R}_{red}(\mathbf{x},\alpha)=\frac{1}{2}\mathbf{x}^{T}(\mathbf{x}-f(\mathbf{x},\alpha))$, Equation 60 implies that an ideal and locally homogeneous denoiser has the form $f(\mathbf{x})=\nabla f(\mathbf{x})\mathbf{x}$, which means that under these conditions, the RED regularization can be thought of as a (pseudo-quadratic) energy function:
@@
An alternative approach to solving inverse problems involves generating samples from the posterior distribution. In other words, given measurements $\mathbf{y}$, we aim to generate samples $\mathbf{x}$ that follow the distribution $P(\mathbf{x}|\mathbf{y})$. We can achieve posterior sampling from $P(\mathbf{x}|\mathbf{y})$ by modifying equation (49) as follows:
@@
Denoising diffusion models are rapidly gaining attention as a powerful tool for solving inverse problems in various fields. This is often achieved by combining the strengths of diffusion models with additional approximations or techniques, as evidenced by a growing body of research (e.g., [63, 47, 64, 65, 66, 67, 68, 69, 70, 71]).

### v2

@@ Abstract
This paper aims to address this gap. We present a clarifying perspective on denoisers, their structure, and desired properties. We emphasize the increasing importance of denoising and showcase its evolution into an essential building block for complex tasks in imaging, inverse problems, and machine learning. Despite its long history, the community continues to uncover unexpected and groundbreaking uses for denoising, further solidifying its place as a cornerstone of scientific and engineering practice.
@@ 1 Introduction
The conservation Property 2 guarantees that the ideal denoiser is the gradient of a scalar field. When the Hessian of the scalar field is bounded, this also implies that $f(\mathbf{x},\alpha)$ is a Lipschitz map with some constant $M(\alpha)$:
@@
We naturally expect $f(\mathbf{0},\alpha)=\mathbf{0}$ for all $\alpha$; therefore, this Lipschitz condition implies $\|f(\mathbf{x},\alpha)\|\leq M(\alpha)\|\mathbf{x}\|$. A non-expansive denoiser would require that $M(\alpha)\leq 1$. In the statistics literature, such operators are called shrinking smoothers [3, 4].
@@ 3.1 Bayesian Denoisers
where $\phi(\mathbf{u})=-\log P(\mathbf{u})$ is the negative log-prior on the space of ”clean” images, and $\alpha$ is proportional to the noise variance. It would appear that the MAP denoiser does not have a closed form. However, the expression (17) is also known in the optimization literature as a proximal operator [21] when $\phi$ is convex, quasi-convex, or a difference of convex functions. It is well-known [22] that to every proximal operator $f$ there corresponds a convex scalar-valued function $\psi$ such that $f=\nabla\psi$.
@@
where $\tilde{\phi}_{\alpha}(\mathbf{x})=-\log P(\mathbf{x},\alpha)$. This is more or less identical to the form of the MAP denoiser in (18). Figure 4 illustrates the MMSE denoiser for the scalar case with $L_{1}$ penalization, showcasing its behavior across various $\alpha$ values. A comparison between the MMSE and Maximum A Posteriori (MAP) estimators is presented in Figure 5.
@@
Data-driven MMSE denoisers: The typical modern supervised approach to image denoising is to train a deep neural network with pairs of clean and noisy images, where the noise is often modeled as additive white and Gaussian(AWGN). Let’s assume we have image pairs $(\mathbf{u},\mathbf{x})\sim P(\mathbf{u},\mathbf{x})$ where $\mathbf{u}$ represents a clean image, and $\mathbf{x}$ is the noisy observation obtained by adding AWGN with a known standard deviation to $\mathbf{u}$: $\mathbf{x}=\mathbf{u}+\mathbf{e}$, where $\mathbf{e}\sim\mathcal{N}(0,\sigma^{2})$.
@@
As mentioned earlier, the MMSE denoiser is the average of all plausible clean signals given the noisy observation. This averaging can lead to a loss of details and a blurry appearance, especially when the noise level is high.
Denoising Autoencoders (DAEs) are a prime example of data-driven MMSE denoisers [27]. These neural networks excel at learning robust data representations by training on noisy input and striving to reconstruct the original, clean data. This makes them not only valuable for denoising but also for tasks like data compression and feature extraction.
@@ 3.2.1 Approximation of Energy-based Denoisers
Though these facts were neither historically clarified, nor the original motivation for their development [29], denoisers of the form (31) have always been heuristic/empirical approximations to the MMSE. These denoisers were hugely popular and effective (e.g. [32, 11]) for decades before the more recent introduction of neural networks. More recent work by Scarvelis et al. [33] explores the use of a specific kernel approach to create a ”closed-form” diffusion model that operates directly on the training set, without the need for training a neural network.
@@ 4 Denoising, the Score Function, and Generative Modeling
At first blush, it is not at all clear why this is a reasonable procedure. Yet there are a number of ways [45, 46, 44, 47, 48] to motivate this idea -perhaps none more direct than by using Tweedie’s formula introduced earlier in Eq. (20).
@@
Despite its elegance, the MMSE estimator is typically difficult to compute, or entirely inaccessible. Therefore as a proxy, often other denoisers are used, which may only be rough approximations of the MMSE (Eq. (43)).
@@ 4.1 Denoising as the Engine of Diffusion Models
Denoising Diffusion Models (DDMs) [46, 44, 51, 52, 48] have become an important area of research in generative modeling. They operate by progressively corrupting training data with noise until it’s indistinguishable from random noise, then learning to systematically reverse this corruption. By training a model to iteratively denoise, it gains the ability to generate entirely new, coherent data samples from a starting point of pure noise, effectively converting noise into meaningful structures like images or other data forms (Figure 6).
@@
which we call a residual flow. As we’ve described in the previous sections, lack of access to the MMSE denoiser forces us to select a different denoiser and therefore solve only an approximate version of the desired Eq. (49).
@@ 5 Denoisers in the Context of Inverse Problems
Over the last several decades, a vast number of choices for the regularizer $\mathbf{R}(\mathbf{x},\alpha)$ have been proposed with varying degrees of success. Early approaches often relied on hand-designed priors to encourage desired properties in the solution, such as sparsity or smoothness [56, 57, 58, 59]. Iterative Shrinkage/Thresholding (IST) algorithms [60, 57, 61, 62, 63] utilize the shrinkage/thresholding function (Moreau proximal mapping) derived from the regularizer $\mathbf{R}$ [62] to solve optimization problems. However, the non-smoothness of many regularizers and the scale of these problems pose computational challenges. Proximal methods like FISTA [64] and ADMM [65, 66] present more efficient solutions by leveraging the proximal operator, which can be interpreted as applying a denoising step to intermediate solutions.
More recently, and independently of the machine learning literature, a fascinating connection has emerged between denoising algorithms and inverse problems. Powerful denoising algorithms, particularly those leveraging deep learning, have been shown to implicitly encode strong priors about natural signals. By incorporating these denoisers into the optimization framework, we can effectively leverage their learned priors to achieve state-of-the-art performance in various inverse problems  [67, 68, 69, 70, 71, 37, 72, 73]. This approach effectively blurs the lines between traditional regularization techniques and modern denoising methods, offering a new paradigm for solving inverse problems.
@@
for the gradient in (57), we have the following expression for the gradient of the objective:
@@
In the particular case of RED: $\mathbf{R}_{red}(\mathbf{x},\alpha)=\frac{1}{2}\mathbf{x}^{T}(\mathbf{x}-f(\mathbf{x},\alpha))$, Equation (60) implies that an ideal and locally homogeneous denoiser has the form $f(\mathbf{x})=\nabla f(\mathbf{x})\mathbf{x}$, which means that under these conditions, the RED regularization can be thought of as a (pseudo-quadratic) energy function:
@@
An alternative approach to solving inverse problems involves generating samples from the posterior distribution using a pretrained prior in the form of a denoiser [80, 81, 82]. In other words, given measurements $\mathbf{y}$, we aim to generate samples $\mathbf{x}$ that follow the distribution $P(\mathbf{x}|\mathbf{y})$. We can achieve posterior sampling from $P(\mathbf{x}|\mathbf{y})$ by modifying equation (49) as follows:
@@
Denoising diffusion models are rapidly gaining attention as a powerful tool for solving inverse problems in various fields. This is often achieved by combining the strengths of diffusion models with additional approximations or techniques, as evidenced by a growing body of research (e.g., [83, 48, 84, 80, 81, 85, 86, 87, 88, 89, 90]).
@@ 6 Conclusions
Acknowledgments
The authors extend their sincere gratitude to Mojtaba Ardakani, Mario A. T. Figueiredo, Ulugbek Kamilov, José Lezama, Ian Manchester and Miki Rubinstein for their valuable feedback.
We also extend our sincere thanks to the vast research community whose dedication over decades has driven remarkable progress in denoising. The advancements in this field are a testament to collective effort, and it is beyond the scope of this work to fully acknowledge the extensive and diverse body of literature on this topic.